Assume the half-life of the first order decay of radioactive isotope takes about 1
year (365 days). How long will it take the radioactivity of that isotope to decay by
60%?
1
Expert's answer
2020-10-14T14:03:38-0400

For first order reaction ,we known

Half life time (t1/2)=0.693/k

Or ,k=(0.693/365)=0.00189 days^-1

For first order reaction

t=(1/0.00189)ln(100/60)=270.27

Or, t=270days,hence 270 days required